OpenStudy (idealist10):

Find all solutions of (x^2)(y)(y')=(y^2-1)^(3/2).

OpenStudy (idealist10):

@ganeshie8

ganeshie8 (ganeshie8):

separable

OpenStudy (idealist10):

I got (x^2)*dy/dx=(y^2-1)^(3/2) divided by y, but what's next?

ganeshie8 (ganeshie8):

next separate the variables : arrange all the y terms on dy side and x terms on dx side

ganeshie8 (ganeshie8):

\[\large \begin{align}\\ x^2y\dfrac{dy}{dx}& =(y^2-1)^{3/2} \end{align}\]

ganeshie8 (ganeshie8):

\[\large \begin{align}\\ x^2y\dfrac{dy}{dx}& =(y^2-1)^{3/2} \\~\\ \dfrac{ydy}{(y^2-1)^{3/2} }& =\dfrac{dx}{x^2}\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

integrate ^^

OpenStudy (idealist10):

How do I integrate y*dy/(y^2-1)^(3/2)?

OpenStudy (idealist10):

u-substitution?

ganeshie8 (ganeshie8):

yes :)

OpenStudy (idealist10):

So I got -1/(sqrt(y^2-1))=-1/x+C, how do I solve for y?

ganeshie8 (ganeshie8):

you may leave the answer in implicit form, don't need to solve it explicitly unless you're asked to do so

OpenStudy (idealist10):

I got sqrt(y^2-1)=x-1/c, am I right?

ganeshie8 (ganeshie8):

1/(sqrt(y^2-1))=1/x+C maybe leave it like this

ganeshie8 (ganeshie8):

if you flip it, you would get : sqrt(y^2-1) = x/(1+Cx)

OpenStudy (idealist10):

Thank you!